Find the particular solution of the following differential equation using the indeterminate coefficients method. Also find the general solution.

{y}''+4{y}'+3y=cosx
Answer #1


у+чэ +3) Soletion Given fliad у+++2)**Я COSX оw we go to . 9 +421 - 23 - Чт +3 = 0 = m* + 2 + 2 0 2) ww(m+3) 19 (12) =0 ( Yp= A cosa + B sima = A (-Sime) + B cosa) y es YpCx) = - A simz + Bcosy y oce) -A (cosa) + BC-Sima) ㅋ - A cosa -B sind. =) y®B value in ea 0 2A +49 B = 2A +4() =1 43=1 2 At ir 4 2 2A = Jln 7) 2A = s 2 alh ک y cz ACOSX + B Sint 2 5 ypcz) 3 cosa te s